# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 3.3 Lax equations
To derive Lax equations for the Lax matrix (3.56) of the cyclotomic Gaudin model, let us return to the hierarchy of Lax equations (2.7) induced by the family of Hamiltonians in involution with respect to {,}$\{\,\,,\,\}_{R}$. In our current setup, {,}$\{\,\,,\,\}_{R}$ is the Lie-Poisson bracket on $\widetilde{\mathfrak{g}}_{D}^{(1)}$ associated with the linear map $R=P_{+}-P_{-}$. Invariant functions on $\widetilde{\mathfrak{g}}_{D}^{(1)}$ take the form
$H_{p,r}\colon Y\in\widetilde{\mathfrak{g}}_{D}^{(1)}\longmapsto\operatorname{Res}_{\lambda_{r}=0}\big{(}\ell_{p,r}(\lambda_{r})\operatorname{Tr}\big{(}Y_{r}(\lambda_{r})^{p+1}\big{)}\big{)}d\lambda,\quad p\geq 1,\quad r\in S\,,$ (3.57)
where $\ell_{p,r}(\lambda_{r})\in\mathbb{C}((\lambda_{r}))$ is a collection of Laurent polynomials for $p\geq 1$ and $r\in S$. It follows from Proposition 3.1 that for these functions to be non-trivial, the Laurent polynomials $\ell_{p,r}(\lambda_{r})$ for $r\in\{0,\infty\}$ should be chosen such that $\ell_{p,r}(\lambda_{r})Y_{r}(\lambda_{r})^{p}\in\widetilde{\mathfrak{g}}_{r}^{\Gamma,0}$, while $\ell_{p,r}(\lambda_{r})$ for $r\in\{1,\ldots,N\}$ can be any Laurent polynomials. Let us then choose
$\ell_{p,r}(\lambda_{r})=\iota_{\lambda_{r}}\frac{\lambda^{p}}{p+1}\quad\text{for}\,\,r\in\{0,\infty\}\quad\text{and}\quad\ell_{p,r}(\lambda_{r})=\iota_{\lambda_{r}}\frac{T\lambda^{p}}{p+1}\quad\text{for}\,\,r\in\{1,\ldots,N\}\,.$ (3.58)
@@
$\partial_{t_{p}^{r}}\iota_{\lambda}L=\big{[}R_{\pm}\nabla H_{p,r}(\iota_{\lambda}L),\iota_{\lambda}L\big{]}\,.$ (3.59)
@@
$\lim_{\epsilon\rightarrow 0}\frac{H_{p,r}(\iota_{\lambda}L+\epsilon\eta)-H_{p,r}(\iota_{\lambda}L)}{\epsilon}=\big{\langle}\eta,\nabla H_{p,r}(\iota_{\lambda}L)\big{\rangle}\,,$ (3.60)
for all $\eta\in\widetilde{\mathfrak{g}}_{D}^{(1)}$. Using Proposition 3.2, we may decompose this gradient as follows
$\nabla H_{p,r}(\iota_{\lambda}L)=N_{r}^{(p)}+\iota_{\lambda}h_{r}^{(p)}\,,\quad N_{r}^{(p)}\in\widetilde{\mathfrak{g}}_{D+}^{(0)}\,,\quad h_{r}^{(p)}\in\mathcal{F}_{D}^{\Gamma}\,,$ (3.61)
@@
$\partial_{t_{p}^{r}}\iota_{\lambda}L=\big{[}R_{-}\nabla H_{p,r}(\iota_{\lambda}L),\iota_{\lambda}L\big{]}=-\big{[}P_{-}\big{(}\nabla H_{p,r}(\iota_{\lambda}L)\big{)},\iota_{\lambda}L\big{]}=-\big{[}\iota_{\lambda}h_{r}^{(p)},\iota_{\lambda}L\big{]}\,.$ (3.62)
Since $\iota_{\lambda}$ is an embedding which also clearly commutes with the Lie brackets $[\cdot,\cdot]:\mathcal{F}_{D}^{\Gamma}\times\Omega_{D}^{\Gamma}\to\Omega_{D}^{\Gamma}$ and $[\cdot,\cdot]:\widetilde{\mathfrak{g}}_{D}^{(0)}\times\widetilde{\mathfrak{g}}_{D}^{(1)}\to\widetilde{\mathfrak{g}}_{D}^{(1)}$, the above equation implies
$\partial_{t_{p}^{r}}L=-\big{[}h_{r}^{(p)},L\big{]}\,.$ (3.63)
In order to calculate $h_{r}^{(p)}\in\mathcal{F}_{D}^{\Gamma}$, by virtue of Proposition 3.4 it is sufficient to restrict $\eta$ in (3.60) to live in $\widetilde{\mathfrak{g}}_{D+}^{(1)}$. We then have
$\big{\langle}\eta,N_{r}^{(p)}\big{\rangle}=0\quad\text{for all}\,\,\eta\in\widetilde{\mathfrak{g}}_{D+}^{(1)}\,,\,\,N_{r}^{(p)}\in\widetilde{\mathfrak{g}}_{D+}^{(0)}\,,$ (3.64)
@@
$T\sum_{s=1}^{N}\operatorname{Res}_{\lambda_{s}=0}\operatorname{Tr}\big{(}\eta_{s}(\lambda_{s})\iota_{\lambda_{s}}h_{r}^{(p)}\big{)}+\sum_{s\in\{0,\infty\}}\operatorname{Res}_{\lambda_{s}=0}\operatorname{Tr}\big{(}\eta_{s}(\lambda_{s})\iota_{\lambda_{s}}h_{r}^{(p)}\big{)}\,,$
@@
$\lim_{\epsilon\rightarrow 0}\frac{H_{p,r}(\iota_{\lambda}L+\epsilon\eta)-H_{p,r}(\iota_{\lambda}L)}{\epsilon}=(p+1)\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}\big{(}\ell_{p,r}(\lambda_{r})\eta_{r}(\lambda_{r})\iota_{\lambda_{r}}L^{p}\big{)}\,,\quad r\in S\,,$ (3.65)
for any $\eta_{s}(\lambda_{s})\in\widetilde{\mathfrak{g}}_{s+}$, $s\in\{1,\ldots,N\}$, and $\eta_{s}(\lambda_{s})\in\widetilde{\mathfrak{g}}_{s+}^{\Gamma,1}$, $s\in\{0,\infty\}$. By definition of $\ell_{p,r}(\lambda_{r})$ in (3.58), this implies
$\big{(}\iota_{\lambda_{s}}h_{r}^{(p)}\big{)}_{-}=\begin{dcases}0&\text{for}\,\,s\neq r\\ \big{(}\iota_{\lambda_{r}}\lambda^{p}\iota_{\lambda_{r}}L^{p}\big{)}_{-}&\text{for}\,\,s=r\,,\end{dcases}$ (3.66)
@@
$\partial_{t_{1}^{r}}L=-\big{[}h_{r}^{(1)},L\big{]}\quad\text{with}\quad h_{r}^{(1)}=\begin{dcases}\dfrac{A_{0}^{(1)}}{\lambda}&\text{for}\,\,r=0\\ \dfrac{1}{T}\sum_{k=0}^{T-1}\dfrac{\omega^{k}\zeta_{r}\sigma^{k}A_{r}}{\lambda-\omega^{k}\zeta_{r}}&\text{for}\,\,r\in\{1,\ldots,N\}\,.\end{dcases}$ (3.67)
Finally, we turn to the primary goal of this paper which is to give a Lagrangian multiform for the cyclotomic Gaudin model that will provide a variational description of the hierarchy of Lax equations in (3.63). Having just described the cyclotomic Gaudin model within the Lie dialgebra framework, we now have all the necessary ingredients to achieve this goal: the non-dynamical element $\Lambda\in\Omega_{D}^{\Gamma}$ that fixes the phase space, the field element $\varphi_{+}\in\widetilde{G}_{D+}^{(0)}$ containing the dynamical degrees of freedom, the linear map $R$ that equips the phase space with the required Poisson structure, and the invariant functions $H_{p,r}$ which induce non-trivial equations of motion with respect to this Poisson structure.
@@ 3.4 Lagrangian multiform for the cyclotomic Gaudin hierarchy
$\mathscr{L}=\sum_{p=1}^{M}\sum_{r\in S\setminus\{\infty\}}\mathscr{L}_{p,r}\mathrm{d}t_{p}^{r}$ (3.68)
using the expression (2.10) for the Lagrangian coefficients. The elementary times $t_{k}$ that appear in (2.10) are now naturally labelled by a pair of indices $p\geq 1$ and $r\in S$, namely we now have elementary times $t_{p}^{r}$, associated with the corresponding Hamiltonians (3.57). Explicitly,
@@
$\mathscr{L}_{p,r}=\sum_{s=1}^{N}\operatorname{Tr}\big{(}A_{s}\partial_{t_{p}^{r}}\phi_{s}\phi_{s}^{-1}\big{)}+\operatorname{Tr}\big{(}A_{0}^{(0)}\partial_{t_{p}^{r}}\phi_{0}^{(0)}\phi_{0}^{(0)\,-1}\big{)}-H_{p,r}(\iota_{\lambda}L)\,,$ (3.69)
@@
$H_{p,r}(\iota_{\lambda}L)=\operatorname{Res}_{\lambda_{r}=0}\big{(}\ell_{p,r}(\lambda_{r})\operatorname{Tr}(\iota_{\lambda_{r}}L^{p+1})\big{)}d\lambda,\quad r\in\{0,1,\ldots,N\}\,,$ (3.70)
@@
$\ell_{p,0}=\iota_{\lambda_{0}}\frac{\lambda^{p}}{p+1}\quad\text{and}\quad\ell_{p,r}=\iota_{\lambda_{r}}\frac{T\lambda^{p}}{p+1}\quad\text{for}\,\,r\in\{1,\ldots,N\}\,.$
The Lagrangian $1$-form $\mathscr{L}$ satisfies the corner equations (1.4)-(1.5) of the multi-time Euler-Lagrange equations, while the standard Euler-Lagrange equations for $\mathscr{L}_{p,r}$ give the hierarchy of Lax equations in (3.63). Further, on solutions of (3.63), we have the closure relation
$\partial_{t_{q}^{s}}\mathscr{L}_{p,r}-\partial_{t_{p}^{r}}\mathscr{L}_{q,s}=0\,,$ (3.71)
@@
Let us start by reinterpreting the formula in (2.10) in the present context of the cyclotomic Gaudin model. First, note that on the coadjoint orbit $\mathcal{O}_{\Lambda}$, where the Lagrangian $1$-form $\mathscr{L}$ lives, the role of the Lax matrix is played by the image of $L$ in $\widetilde{\mathfrak{g}}_{D}^{(1)}$ under the embedding $\iota_{\lambda}$, given by (3.2). Next, the bilinear pairing used to define the kinetic part is the one constructed in Proposition 3.1. Furthermore, in the Adler-Kostant-Symes scheme where (locally) $\mathbb{G}_{R}\simeq\mathbb{G}_{+}\times\mathbb{G}_{-}$, we explicitly have $\partial_{t_{p}^{r}}\varphi\cdot_{R}\varphi^{-1}=\partial_{t_{p}^{r}}\varphi_{+}\varphi_{+}^{-1}+\partial_{t_{p}^{r}}\varphi_{-}\varphi_{-}^{-1}$, and since $\iota_{\lambda}L\in V_{-}=\iota_{\lambda}\Omega_{D}^{\Gamma}$, it follows from Proposition 3.4 that only the $\partial_{t_{p}^{r}}\varphi_{+}\varphi_{+}^{-1}$ piece contributes to the kinetic term. Finally, the potential part is simply the restriction of the invariant functions $H_{p,r}$ in (3.57) to $\iota_{\lambda}L$. Therefore, in the present setup, the Lagrangian coefficients in (2.10) can be expressed as
@@
$\mathscr{L}_{p,r}$ $=\big{\langle}\,\iota_{\lambda}L,\partial_{t_{p}^{r}}\varphi_{+}\varphi_{+}^{-1}\,\big{\rangle}-H_{p,r}(\iota_{\lambda}L)$
$=\big{\langle}\,\iota_{\lambda}\Lambda,\varphi_{+}^{-1}\partial_{t_{p}^{r}}\varphi_{+}\big{\rangle}-H_{p,r}(\iota_{\lambda}L),\quad r\in\{0,1,\ldots,N\}$ (3.72)
where in the second step we used the fact (3.2) that $\iota_{\lambda}L=\mathcal{P}_{-}\big{(}\varphi_{+}(\iota_{\lambda}\Lambda)\varphi_{+}^{-1}\big{)}$. The kinetic term can be written out more explicitly in terms of the component fields (3.47) as
@@
$\big{\langle}\,\iota_{\lambda}\Lambda,\varphi_{+}^{-1}\partial_{t_{p}^{r}}\varphi_{+}\,\big{\rangle}$ $=T\sum_{s=1}^{N}\operatorname{Res}_{\lambda_{s}=0}\operatorname{Tr}\big{(}\iota_{\lambda_{s}}\Lambda\,\varphi_{s+}(\lambda_{s})^{-1}\,\partial_{t_{p}^{r}}\varphi_{s+}(\lambda_{s})\big{)}d\lambda$
@@
$\quad+\sum_{s\in\{0,\infty\}}\operatorname{Res}_{\lambda_{s}=0}\operatorname{Tr}\big{(}\iota_{\lambda_{s}}\Lambda\,\varphi_{s+}(\lambda_{s})^{-1}\,\partial_{t_{p}^{r}}\varphi_{s+}(\lambda_{s})\big{)}d\lambda$
$=\sum_{s=1}^{N}\operatorname{Tr}\big{(}A_{s}\partial_{t_{p}^{r}}\phi_{s}\phi_{s}^{-1}\big{)}+\operatorname{Tr}\big{(}A_{0}^{(0)}\partial_{t_{p}^{r}}\phi_{0}^{(0)}\phi_{0}^{(0)\,-1}\big{)}$
@@
$\quad+\partial_{t_{p}^{r}}\operatorname{Tr}\big{(}\phi_{0}^{(1)}\phi_{0}^{(0)\,-1}A_{0}^{(1)}\big{)}+\frac{1}{2}\partial_{t_{p}^{r}}\operatorname{Tr}\big{(}A_{\infty}(\phi_{\infty}^{(1)})^{2}\big{)}\,.$ (3.73)
@@
The last two terms associated to the poles at the origin and at infinity that are total derivatives do not contribute to the Euler-Lagrange equations and therefore can be dropped, leaving us with the required expression for the Lagrangian coefficients.
@@
To close this section, let us present the explicit expressions for the first set of Lagrangian coefficients $\mathscr{L}_{1,r}$ with $r\in\{0,1,\ldots,N\}$. The kinetic terms are obtained by simply substituting $p=1$ in the kinetic part in (3.69), while the potential terms defined by (3.70) read
@@
Upon varying $\mathscr{L}_{1,r}$ with respect to $\phi_{s}$, $s=1,\ldots,N$, $\phi_{0}^{(0)}$, and $\phi_{0}^{(1)}$, it can be checked that the associated Euler-Lagrange equations correspond to the set of Lax equations for $p=1$ in (3.67), as it should be from Theorem 3.6.

### v2

@@ 3.3 Lax equations
To derive Lax equations for the Lax matrix (3.28) of the cyclotomic Gaudin model, let us return to the hierarchy of Lax equations (2.6) induced by the family of Hamiltonians in involution with respect to {,}$\{\,,\,\}_{R}$. In our current setup, {,}$\{\,,\,\}_{R}$ is the Lie–Poisson bracket on $\widetilde{\mathfrak{g}}_{D}^{(1)}$ associated with the linear map $R=P_{+}-P_{-}$. Invariant functions on $\widetilde{\mathfrak{g}}_{D}^{(1)}$ take the form
$H_{p,r}\colon\ Y\in\widetilde{\mathfrak{g}}_{D}^{(1)}\longmapsto\operatorname{Res}_{\lambda_{r}=0}\bigl{(}\ell_{p,r}(\lambda_{r})\operatorname{Tr}\bigl{(}Y_{r}(\lambda_{r})^{p+1}\bigr{)}\bigr{)}{\rm d}\lambda,\qquad p\geq 1,\qquad r\in S,$ (3.29)
where $\ell_{p,r}(\lambda_{r})\in\mathbb{C}((\lambda_{r}))$ is a collection of Laurent polynomials for $p\geq 1$ and $r\in S$. It follows from Proposition 3.1 that for these functions to be non-trivial, the Laurent polynomials $\ell_{p,r}(\lambda_{r})$ for $r\in\{0,\infty\}$ should be chosen such that $\ell_{p,r}(\lambda_{r})Y_{r}(\lambda_{r})^{p}\in\widetilde{\mathfrak{g}}_{r}^{\Gamma,0}$, while $\ell_{p,r}(\lambda_{r})$ for $r\in\{1,\dots,N\}$ can be any Laurent polynomials. Let us then choose
$\ell_{p,r}(\lambda_{r})=\iota_{\lambda_{r}}\frac{\lambda^{p}}{p+1}\qquad\text{for}\quad r\in\{0,\infty\},$
$\ell_{p,r}(\lambda_{r})=\iota_{\lambda_{r}}\frac{T\lambda^{p}}{p+1}\qquad\text{for}\quad r\in\{1,\dots,N\}.$ (3.30)
@@
$\partial_{t_{p}^{r}}\iota_{\lambda}L=\bigl{[}R_{\pm}\nabla H_{p,r}(\iota_{\lambda}L),\iota_{\lambda}L\bigr{]}.$ (3.31)
@@
$\lim_{\epsilon\rightarrow 0}\frac{H_{p,r}(\iota_{\lambda}L+\epsilon\eta)-H_{p,r}(\iota_{\lambda}L)}{\epsilon}=\langle\eta,\nabla H_{p,r}(\iota_{\lambda}L)\rangle$ (3.32)
for all $\eta\in\widetilde{\mathfrak{g}}_{D}^{(1)}$. Using Proposition 3.2, we may decompose this gradient as follows:
$\nabla H_{p,r}(\iota_{\lambda}L)=N_{r}^{(p)}+\iota_{\lambda}h_{r}^{(p)},\qquad N_{r}^{(p)}\in\widetilde{\mathfrak{g}}_{D+}^{(0)},\qquad h_{r}^{(p)}\in\mathcal{F}_{D^{\prime}}^{\Gamma},$
@@
$\partial_{t_{p}^{r}}\iota_{\lambda}L=[R_{-}\nabla H_{p,r}(\iota_{\lambda}L),\iota_{\lambda}L]=-[P_{-}(\nabla H_{p,r}(\iota_{\lambda}L)),\iota_{\lambda}L]=-\bigl{[}\iota_{\lambda}h_{r}^{(p)},\iota_{\lambda}L\bigr{]}.$
Since $\iota_{\lambda}$ is an embedding which also clearly commutes with the Lie brackets $[\cdot,\cdot]\colon\mathcal{F}_{D^{\prime}}^{\Gamma}\times\Omega_{D^{\prime}}^{\Gamma}\to\Omega_{D^{\prime}}^{\Gamma}$ and $[\cdot,\cdot]\colon\widetilde{\mathfrak{g}}_{D}^{(0)}\times\widetilde{\mathfrak{g}}_{D}^{(1)}\to\widetilde{\mathfrak{g}}_{D}^{(1)}$, the above equation implies
$\partial_{t_{p}^{r}}L=-\bigl{[}h_{r}^{(p)},L\bigr{]}.$ (3.33)
In order to calculate $h_{r}^{(p)}\in\mathcal{F}_{D^{\prime}}^{\Gamma}$, by virtue of Proposition 3.4, it is sufficient to restrict $\eta$ in (3.32) to live in $\widetilde{\mathfrak{g}}_{D+}^{(1)}$. We then have
$\bigl{\langle}\eta,N_{r}^{(p)}\bigr{\rangle}=0\qquad\text{for all}\quad\eta\in\widetilde{\mathfrak{g}}_{D+}^{(1)},\quad N_{r}^{(p)}\in\widetilde{\mathfrak{g}}_{D+}^{(0)},$
@@
$T\sum_{s=1}^{N}\operatorname{Res}_{\lambda_{s}=0}\operatorname{Tr}\bigl{(}\eta_{s}(\lambda_{s})\iota_{\lambda_{s}}h_{r}^{(p)}\bigr{)}+\sum_{s\in\{0,\infty\}}\operatorname{Res}_{\lambda_{s}=0}\operatorname{Tr}\bigl{(}\eta_{s}(\lambda_{s})\iota_{\lambda_{s}}h_{r}^{(p)}\bigr{)},$
@@
$\lim_{\epsilon\rightarrow 0}\frac{H_{p,r}(\iota_{\lambda}L+\epsilon\eta)-H_{p,r}(\iota_{\lambda}L)}{\epsilon}=(p+1)\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}\bigl{(}\ell_{p,r}(\lambda_{r})\eta_{r}(\lambda_{r})\iota_{\lambda_{r}}L^{p}\bigr{)},\qquad r\in S,$
for any $\eta_{s}(\lambda_{s})\in\widetilde{\mathfrak{g}}_{s+}$, $s\in\{1,\dots,N\}$, and $\eta_{s}(\lambda_{s})\in\widetilde{\mathfrak{g}}_{s+}^{\Gamma,1}$, $s\in\{0,\infty\}$. By definition of $\ell_{p,r}(\lambda_{r})$ in (3.30), this implies
$\bigl{(}\iota_{\lambda_{s}}h_{r}^{(p)}\bigr{)}_{-}=\begin{dcases}0&\text{for}\ s\neq r,\\ \bigl{(}\iota_{\lambda_{r}}\lambda^{p}\iota_{\lambda_{r}}L^{p}\bigr{)}_{-}&\text{for}\ s=r\end{dcases}$
@@
$\partial_{t_{1}^{r}}L=-\bigl{[}h_{r}^{(1)},L\bigr{]}\qquad\text{with}\quad h_{r}^{(1)}=\begin{dcases}\dfrac{A_{0}^{(1)}}{\lambda}&\text{for}\ r=0,\\ \dfrac{1}{T}\sum_{k=0}^{T-1}\dfrac{\omega^{k}\zeta_{r}\sigma^{k}A_{r}}{\lambda-\omega^{k}\zeta_{r}}&\text{for}\ r\in\{1,\dots,N\}.\end{dcases}$ (3.34)
Finally, we turn to the primary goal of this paper which is to give a Lagrangian multiform for the cyclotomic Gaudin model that will provide a variational description of the hierarchy of Lax equations in (3.33). Having just described the cyclotomic Gaudin model within the Lie dialgebra framework, we now have all the necessary ingredients to achieve this goal: the non-dynamical element $\Lambda\in\Omega_{D^{\prime}}^{\Gamma}$ that fixes the phase space, the field element $\varphi_{+}\in\widetilde{G}_{D+}^{(0)}$ containing the dynamical degrees of freedom, the linear map $R$ that equips the phase space with the required Poisson structure, and the invariant functions $H_{p,r}$ which induce non-trivial equations of motion with respect to this Poisson structure.
@@ 3.4 Lagrangian multiform for the cyclotomic Gaudin hierarchy
$\mathscr{L}=\sum_{p=1}^{M}\sum_{r\in S\setminus\{\infty\}}\mathscr{L}_{p,r}\,\mathrm{d}t_{p}^{r},$
using the expression (2.8) for the Lagrangian coefficients. It will be useful to recall the notations in (3.27) associated with the parameterisation of the cyclotomic Gaudin Lax matrix. The elementary times $t_{k}$ that appear in (2.8) are now naturally labelled by a pair of indices ${p\geq 1}$ and ${r\in S}$, namely we now have elementary times $t_{p}^{r}$, associated with the corresponding Hamiltonians (3.29). Explicitly, we have the following.
@@
$\mathscr{L}_{p,r}=\sum_{s=1}^{N}\operatorname{Tr}\bigl{(}A_{s}\partial_{t_{p}^{r}}\phi_{s}\phi_{s}^{-1}\bigr{)}+\operatorname{Tr}\bigl{(}A_{0}^{(0)}\partial_{t_{p}^{r}}\phi_{0}^{(0)}\phi_{0}^{(0)\,-1}\bigr{)}-H_{p,r}(\iota_{\lambda}L),$ (3.35)
@@
$H_{p,r}(\iota_{\lambda}L)=\operatorname{Res}_{\lambda_{r}=0}\bigl{(}\ell_{p,r}(\lambda_{r})\operatorname{Tr}\bigl{(}\iota_{\lambda_{r}}L^{p+1}\bigr{)}\bigr{)}d\lambda,\qquad r\in\{0,1,\dots,N\},$ (3.36)
@@
$\ell_{p,0}=\iota_{\lambda_{0}}\frac{\lambda^{p}}{p+1}\qquad\text{and}\qquad\ell_{p,r}=\iota_{\lambda_{r}}\frac{T\lambda^{p}}{p+1}\qquad\text{for}\quad r\in\{1,\dots,N\}.$
The Lagrangian $1$-form $\mathscr{L}$ satisfies the corner equations (1.2)–(1.3) of the multi-time Euler–Lagrange equations, while the standard Euler–Lagrange equations for $\mathscr{L}_{p,r}$ give the hierarchy of Lax equations in (3.33). Further, on solutions of (3.33), we have the closure relation
$\partial_{t_{q}^{s}}\mathscr{L}_{p,r}-\partial_{t_{p}^{r}}\mathscr{L}_{q,s}=0$
@@
Let us start by reinterpreting the formula in (2.8) in the present context of the cyclotomic Gaudin model. First, note that on the coadjoint orbit $\mathcal{O}_{\Lambda}$, where the Lagrangian $1$-form $\mathscr{L}$ lives, the role of the Lax matrix is played by the image of $L$ in $\widetilde{\mathfrak{g}}_{D}^{(1)}$ under the embedding $\iota_{\lambda}$, given by (3.26). Next, the bilinear pairing used to define the kinetic part is the one constructed in Proposition 3.1. Furthermore, in the Adler–Kostant–Symes scheme where (locally) $\mathbb{G}_{R}\simeq\mathbb{G}_{+}\times\mathbb{G}_{-}$, we explicitly have $\partial_{t_{p}^{r}}\varphi\cdot_{R}\varphi^{-1}=\partial_{t_{p}^{r}}\varphi_{+}\varphi_{+}^{-1}+\partial_{t_{p}^{r}}\varphi_{-}\varphi_{-}^{-1}$, and since $\iota_{\lambda}L\in V_{-}=\iota_{\lambda}\Omega_{D^{\prime}}^{\Gamma}$, it follows from Proposition 3.4 that only the $\partial_{t_{p}^{r}}\varphi_{+}\varphi_{+}^{-1}$ piece contributes to the kinetic term. Finally, the potential part is simply the restriction of the invariant functions $H_{p,r}$ in (3.29) to $\iota_{\lambda}L$. Therefore, in the present setup, the Lagrangian coefficients in (2.8) can be expressed as
@@
$\mathscr{L}_{p,r}$ ${}=\bigl{\langle}\iota_{\lambda}L,\partial_{t_{p}^{r}}\varphi_{+}\varphi_{+}^{-1}\bigr{\rangle}-H_{p,r}(\iota_{\lambda}L)$
@@
${}=\bigl{\langle}\iota_{\lambda}\Lambda,\varphi_{+}^{-1}\partial_{t_{p}^{r}}\varphi_{+}\bigr{\rangle}-H_{p,r}(\iota_{\lambda}L),\qquad r\in\{0,1,\dots,N\},$ (3.37)
where in the second step we used the fact (3.26) that $\iota_{\lambda}L=\mathcal{P}_{-}\bigl{(}\varphi_{+}(\iota_{\lambda}\Lambda)\varphi_{+}^{-1}\bigr{)}$. The kinetic term can be written out more explicitly in terms of the component fields (3.24) as
@@
$\bigl{\langle}\iota_{\lambda}\Lambda,\varphi_{+}^{-1}\partial_{t_{p}^{r}}\varphi_{+}\bigr{\rangle}$ ${}=T\sum_{s=1}^{N}\operatorname{Res}_{\lambda_{s}=0}\operatorname{Tr}\bigl{(}\iota_{\lambda_{s}}\Lambda\varphi_{s+}(\lambda_{s})^{-1}\partial_{t_{p}^{r}}\varphi_{s+}(\lambda_{s})\bigr{)}{\rm d}\lambda$
@@
$\quad{}+\sum_{s\in\{0,\infty\}}\operatorname{Res}_{\lambda_{s}=0}\operatorname{Tr}\bigl{(}\iota_{\lambda_{s}}\Lambda\varphi_{s+}(\lambda_{s})^{-1}\partial_{t_{p}^{r}}\varphi_{s+}(\lambda_{s})\bigr{)}{\rm d}\lambda$
${}=\sum_{s=1}^{N}\operatorname{Tr}\bigl{(}A_{s}\partial_{t_{p}^{r}}\phi_{s}\phi_{s}^{-1}\bigr{)}+\operatorname{Tr}\bigl{(}A_{0}^{(0)}\partial_{t_{p}^{r}}\phi_{0}^{(0)}\phi_{0}^{(0)\,-1}\bigr{)}$
$\quad{}+\partial_{t_{p}^{r}}\operatorname{Tr}\bigl{(}\phi_{0}^{(1)}\phi_{0}^{(0)\,-1}A_{0}^{(1)}\bigr{)}+\frac{1}{2}\partial_{t_{p}^{r}}\operatorname{Tr}\bigl{(}A_{\infty}\bigl{(}\phi_{\infty}^{(1)}\bigr{)}^{2}\bigr{)}.$
@@
The last two terms associated to the poles at the origin and at infinity are total derivatives and do not contribute to the multi-time Euler–Lagrange equations since dropping them amounts to changing the Lagrangian $1$-form $\mathscr{L}$ by $\mathrm{d}\operatorname{Tr}\bigl{(}\phi_{0}^{(1)}\phi_{0}^{(0)\,-1}A_{0}^{(1)}+\frac{1}{2}A_{\infty}\bigl{(}\phi_{\infty}^{(1)}\bigr{)}^{2}\bigr{)}$ which is a total horizontal differential. Discarding these two terms, we are left with the required expression for the Lagrangian coefficients.
@@
To close this section, let us present the explicit expressions for the first set of Lagrangian coefficients $\mathscr{L}_{1,r}$ with $r\in\{0,1,\dots,N\}$. The kinetic terms are obtained by simply substituting $p=1$ in the kinetic part in (3.35), while the potential terms defined by (3.36) read
@@
$H_{1,r}=\operatorname{Tr}\bigl{(}A_{0}^{(0)}A_{r}\bigr{)}+\frac{\operatorname{Tr}\bigl{(}A_{0}^{(1)}A_{r}\bigr{)}}{\zeta_{r}}+\frac{1}{2T}\sum_{k=0}^{T-1}\operatorname{Tr}\bigl{(}A_{r}\sigma^{k}A_{r}\bigr{)}$
${}+\frac{1}{T}\sum_{s\neq r}\sum_{k=0}^{T-1}\frac{\operatorname{Tr}\bigl{(}A_{r}\sigma^{k}A_{s}\bigr{)}\zeta_{r}}{\zeta_{r}-\omega^{k}\zeta_{s}}+\operatorname{Tr}(A_{r}A_{\infty})\zeta_{r},\qquad r\in\{1,\dots,N\}.$
Upon varying $\mathscr{L}_{1,r}$ with respect to $\phi_{s}$, $s=1,\dots,N$, $\phi_{0}^{(0)}$, and $\phi_{0}^{(1)}$, it can be checked that the associated Euler–Lagrange equations correspond to the set of Lax equations for $p=1$ in (3.34), as it should be from Theorem 3.6.